OpenStudy (anonymous):

6.Create your own third degree polynomial that when divided by x + 2 has a remainder of –4.

OpenStudy (cwrw238):

x^3 + x^2 -2x - 4 = 0

OpenStudy (cwrw238):

i used the remainder theorem f(-2) = -4

OpenStudy (anonymous):

1x^2-1x – 4/x+2 <<<<< is the answer

OpenStudy (cwrw238):

plug -2 into this gives (-2)^2 + 2 +2 + 2 = 6 that a remainder 6 not -4

OpenStudy (cwrw238):

if a polynomial is divided by (x+2) and the remainder is -4, then if you plug x = -2 into the polynomial then value of polynomial will be -4

OpenStudy (anonymous):

i still dnt understand

OpenStudy (cwrw238):

ok lets suppose we divide x^2 - 4x + 1 by x - 1: x - 3 ------------- x - 1) x^2 - 4x + 1 x^2 - x -------- -3x + 1 -3x + 3 ------- -2 remainder is -2 now replace x in polynomial by 1 ( the 1 fro x-1) we will get -2 (1)^2 - 4(1) + 1 = 1 - 4 _ 1 = -2

OpenStudy (cwrw238):

perhaps you haven't come across the remainder theorem before sassy?

OpenStudy (campbell_st):

try P(x) = (x+2)^2 -4